You own a $36,800 portfolio that is invested in Stocks A and B. The portfolio beta is equal to the market beta. Stock A has an expected return of 15 percent and has a beta of 2. Stock B has a beta of 0.5. What is the value of your investment in Stock A?

Multiple Choice

$10,055


$16,601


$18,539

$12,267
Answer #1

Let investment in A=$x

Hence investment in B=(36800-x)

Portfolio beta=Respective beta*Respective weight

1=(x/36800*2)+(36800-x)/36800*0.5[Beta of market=1;Beta of risk-free assets=0]   

(1*36800)=2x+18400-0.5x

36800=2x+18400-0.5x

x=(36800-18400)/(2-0.5)

=$12,267(Approx)=investment in A